A recipe calls for 1/3 cup of sugar and 2 cups of flour. How much sugar should be used with 7 cups of flour?
Answer by stanbon(75887)

A recipe calls for 1/3 cup of sugar and 2 cups of flour. How much sugar should be used with 7 cups of flour?
--------
Use a proportion::
x/7 = (1/3)/2
-----
x = 7(1/6)
-----
x = 7/6 cup of sugar
====================
Cheers,
Stan H.
====================